OpenStudy (liddoislander):

Solve the following system of equations. x + y = 14 -x + y = 16

OpenStudy (freckles):

Try adding the equations together

OpenStudy (freckles):

x+y=14 +(-x+y=16) ----------- (x+(-x))+(y+y)=14+16

OpenStudy (freckles):

simplify that and then solve for y

OpenStudy (jdoe0001):

hint : \(\begin{array}{rllll} \cancel{ x } + y = 14 \\ \cancel{ -x } + y = 16 \\\hline\\ \square ?+\square ?=\square ? \end{array}\)

OpenStudy (liddoislander):

@bibby can u please help